OpenStudy (across):

What is the smallest positive number that is evenly divisible by all of the numbers from 1 to 20?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

why isn't it \[20!\]?

OpenStudy (anonymous):

oh i see why never mind

OpenStudy (anonymous):

1745944200, though it seems awfully large

hero (hero):

I was thinking about it backwards

OpenStudy (anonymous):

its\[2^4\cdot3^2\cdot5\cdot7\cdot11\cdot13\cdot17\cdot19\]

OpenStudy (saifoo.khan):

\[20! = 2.43290200817664 × 10^18\]

OpenStudy (anonymous):

232792560

OpenStudy (anonymous):

ok you know you need \[2,3,5,7,13,17,19\] as factors, now lets see what powers you need. need \[2^4\] to get 16 need \[3^2\] to get 9, and i think that is it because \[5^2=25\] too big, so i am going to say \[2^4\times 3^2\times 5\times7\times 13\times 17\times 19\]

OpenStudy (anonymous):

dont forget 11 satellite :)

OpenStudy (across):

aaaand joemath hits the jackpot.

OpenStudy (anonymous):

i will let joemath correct me if i am wrong.

OpenStudy (anonymous):

\[2^4\times 3^2\times 5\times7\times 11\times 13\times 17\times 19\]

OpenStudy (anonymous):

forgot one didn't i

OpenStudy (anonymous):

the idea was right though. just look and see what powers of primes are included from 1 to 20.

OpenStudy (anonymous):

oh wow i didn't see you had the answer miles above...

OpenStudy (anonymous):

no one see's my posts =/ lolol jk

OpenStudy (asnaseer):

@across - I thought you said "evenly divisible"? 232792560 / 17 = 14549535 which is not an even number?

OpenStudy (anonymous):

I missed the powers of primes part :(

OpenStudy (anonymous):

by evenly divisible, i understood it as "no remainder"

OpenStudy (anonymous):

That's how I took it

OpenStudy (asnaseer):

sorry I meant 232792560 / 16 = 14549535

OpenStudy (across):

Even divisibility often implies a division leaving no remainder.

OpenStudy (asnaseer):

ok - my misunderstanding then.